OpenStudy (anonymous):

A man 6 ft tall walks at the rate of 5 ft/sec toward a streetlight that is 16 ft tall. At what rate is the length of the shadow changing when he is 10 ft from the base of the light?

OpenStudy (anonymous):

I have a picture drawn out but I don't understand how to set up an equation and take its derivative.

OpenStudy (ddcamp):



OpenStudy (anonymous):

Oh I wasnt using two variables!

OpenStudy (anonymous):

But after I get this what do I do?

OpenStudy (anonymous):

I think I have the right answer, y'(t)= 6/10 *dx/dt y'(t) = 6/10 * 5 y'(t) = 3 ft/sec

OpenStudy (radar):

An excellent drawing which clearly shows the necessary relationships to solve.